OpenStudy (yellowboy):

If f(x) = x2 + x and g(x) = x − 3, find f(g(7)).

OpenStudy (yellowboy):

20 28 32 53

OpenStudy (yellowboy):

53 i got it now

jhonyy9 (jhonyy9):

like a first step g(7) = 7-3 so g(7)=4 and now f(4) = 4^2 +4 = 16+4 =20 hope helped you